Inactive Tutor answered 10/20/19
Tutor
New to Wyzant
There is a 100% chance that an event E will or will not happen. If p(E) = 0.29 = 29% that it WILL happen, then p(^E) = 1-0.29 = 0.71 = 71% that it will not happen.